# MSC1, MSC2, MSC3

### v1

@@ Abstract
Sequences of parametrized Lyapunov equations can be encountered in many application settings. Moreover, solutions of such equations are often intermediate steps of an overall procedure whose main goal is the computation of quantities of the form $f(X)$ where $X$ denotes the solution of a Lyapunov equation. We are interested in addressing problems where the parameter dependency of the coefficient matrix is encoded as a low-rank modification to a seed, fixed matrix. We propose two novel numerical procedures that fully exploit such a common structure. The first one builds upon recycling Krylov techniques, and it is well-suited for small dimensional problems as it makes use of dense numerical linear algebra tools. The second algorithm can instead address large-scale problems by relying on state-of-the-art projection techniques based on the extended Krylov subspace. We test the new algorithms on several problems arising in the study of damped vibrational systems and the analyses of output synchronization problems for multi-agent systems. Our results show that the algorithms we propose are superior to state-of-the-art techniques as they are able to remarkably speed up the computation of accurate solutions.
@@
Novel, efficient methods for solving sequences of parametrized Lyapunov equations have been designed. By taking full advantage of the parameter dependency structure we are able to tackle both small-scale and large-scale problems. The numerical results show the superiority of the proposed methods when compared to state-of-the-art routines, especially in terms of computational time.
@@ 1 Introduction and setting
Here $A_{0},Q\in{\mathbb{R}}^{n\times n}$, $B_{l},B_{r}\in{\mathbb{R}}^{n\times k}$are fixed matrices and the matrix $D(\mathsf{v})=\operatorname{diag}{(v_{1},v_{2},\ldots,v_{k})}$ contains parameters $v_{i}\in{\mathbb{R}}\setminus\left\{0\right\}$, for $i=1\ldots,k$, encoded in the parameter vector $\mathsf{v}=\begin{bmatrix}v_{1},v_{2},\ldots,v_{k}\end{bmatrix}^{\mkern-1.5mu\mathsf{T}}\in{({\mathbb{R}}\setminus\left\{0\right\})}^{k}$ [^1] For the sake of simplicity, we assume the data of our problem to be real. Our algorithms can handle complex data by applying straightforward modifications.. Note that, without loss of generality, we assume that the parameters in $\mathsf{v}$ are nonzero. Indeed, if some parameters were equal to zero, this would be equivalent to a structured update of smaller rank.
@@
We are especially interested in devising efficient procedures for calculating $f(X(\mathsf{v}))$ for a large number of different parameter vectors $\mathsf{v}$, for a given function $f$ that is assumed to be additive, i.e. $f(X+Y)=f(X)+f(Y)$. Indeed, $f(X(\mathsf{v}))$ often corresponds to the target value we are interested in computing, with calculating $X(\mathsf{v})$ only being an intermediate step. For instance, when computing the $H_{2}$-norm of a linear system, one is often interested in computing $f(X(\mathsf{v}))=\operatorname*{trace}(X(\mathsf{v}))$ or $f(X(\mathsf{v}))=\operatorname*{trace}(EX(\mathsf{v}))$ for some matrix $E$. An efficient computation of $f(X(\mathsf{v}))$, and thus the fast solution of (1), are crucial to obtain competitive numerical schemes in terms of both computational performance and memory requirements for the applications mentioned above.
@@
The first procedure we suggest in this paper builds upon the work in [46, 24], which we enhance with a recycling Krylov technique. This routine makes use of dense linear algebra tools so that it is well suited for small dimensional problems, i.e., in the case of moderate values of the matrix dimension $n$. The second routine addresses the large-scale case employing projection methods for linear matrix equations.
@@
In Section 2 we present the main contribution of this work. In particular, a recycling Krylov-like method, which is able to efficiently deal with small dimensional problems, is illustrated in Section 2.1. As already mentioned, this scheme employs dense numerical linear algebra tools, e.g., the full eigenvalue decomposition of $A_{0}$ is required, so that it cannot be applied in the large-scale setting. We address the latter scenario in Section 2.2, where a sophisticated projection technique, based on the extended Krylov subspace (17), is illustrated. Some details about two application settings where equations of the form (1) are encountered are reported in Section 3. In particular, damped vibrational systems are treated in Section 3.1, whereas in Section 3.2 multi-agent systems are considered. A panel of diverse numerical experiments displaying the effectiveness of our methodology is presented in Section 4, while Section 5 collects our conclusions.
@@ 2 Computing $f(X(\mathsf{v}))$
2.1 Small-scale setting: recycling Krylov methods
@@
The strategy proposed in this section builds upon the work in [46, 24]. We extend their approach to the more general class of problems of the form (1), and we allow for more than one parameter. The scheme we are going to present relies on the computation of the full eigenvalue decomposition of $A_{0}$, namely $A_{0}=Q_{0}\Lambda_{0}Q_{0}^{-1}$ with $Q_{0}\in{\mathbb{C}}^{n\times n}$ and $\Lambda_{0}=\operatorname{diag}(\lambda_{1},\ldots,\lambda_{n})$, $\lambda_{i}\in{\mathbb{C}}$, $i=1,\ldots,n$.
@@
We now show how to compute the two terms in the right-hand side of (2.1). The main goal here is to divide the operations that depend on $\mathsf{v}$ from those that do not, as much as possible, such that the $\mathsf{v}$-independent calculations can be performed offline.
@@
While ${\mathbf{D}(\mathsf{v})}^{-1}=\begin{bmatrix}{D(\mathsf{v})}^{-1}\otimes I&\\ &I\otimes{D(\mathsf{v})}^{-1}\\ \end{bmatrix}$ is a diagonal matrix, the structure of $\mathbf{N}^{\mkern-1.5mu\mathsf{T}}{\mathbf{L}}_{0}^{-1}\mathbf{M}$ is more involved. A first study of its structure can be found in [46]. Unlike what is done in [46], here we exploit the diagonal pattern of ${\mathbf{L}}_{0}^{-1}$ which leads to a useful representation of the blocks of $\mathbf{N}^{\mkern-1.5mu\mathsf{T}}{\mathbf{L}}_{0}^{-1}\mathbf{M}$ when the latter is viewed as a $2\times 2$ block matrix. Such a block structure will help us to design efficient preconditioners for the solution of the linear system in (10) as well.
@@
where $\delta_{q,s}$ denotes the Kronecker delta. The relation in (2.1) shows that $\mathbf{N}_{1}^{\mkern-1.5mu\mathsf{T}}{\mathbf{L}}_{0}^{-1}\mathbf{M}_{1}$ is a $k\times k$ block matrix whose blocks are all diagonal. In particular, the block in the $(t,p)$ position is given by matrix $\mathop{\mathrm{diag}}\left({\mathcal{L}}\left((Q_{0}^{\mkern-1.5mu\mathsf{T}}B_{r}e_{t})\mathop{\circ}(Q_{0}^{-1}B_{l}e_{p})\right)\right)$.
@@
Unlike its diagonal blocks, the off-diagonal blocks of $\mathbf{N}^{\mkern-1.5mu\mathsf{T}}{\mathbf{L}}_{0}^{-1}\mathbf{M}$, namely $\mathbf{N}_{2}^{\mkern-1.5mu\mathsf{T}}{\mathbf{L}}_{0}^{-1}\mathbf{M}_{1}$, and $\mathbf{N}_{1}^{\mkern-1.5mu\mathsf{T}}{\mathbf{L}}_{0}^{-1}\mathbf{M}_{2}$, do not possess a structured sparsity pattern. Nevertheless, we can still apply the same strategy we employed above to construct $\mathbf{N}_{2}^{\mkern-1.5mu\mathsf{T}}{\mathbf{L}}_{0}^{-1}\mathbf{M}_{1}$, and $\mathbf{N}_{1}^{\mkern-1.5mu\mathsf{T}}{\mathbf{L}}_{0}^{-1}\mathbf{M}_{2}$ while avoiding the explicit computation of $\mathbf{M}_{1}$, $\mathbf{M}_{2}$, $\mathbf{N}_{1}$, and $\mathbf{N}_{2}$. If $i=(s-1)k+t$ and $j=(p-1)n+q$, we have
@@
We employ the $s_{\ell}$ eigenvectors corresponding to the $s_{\ell}$ eigenvalues of smallest magnitude [^2] This is the default strategy in [34]., but different approximate eigenvectors can be used as well, see [33, Section 2.4].
GCRO-DR allows for restarting and preconditioning. In particular, thanks to the structure of ${\mathbf{D}(\mathsf{v})}^{-1}-\mathbf{N}^{\mkern-1.5mu\mathsf{T}}{\mathbf{L}}_{0}^{-1}\mathbf{M}$ we are able to design efficient preconditioning operators, which can be cheaply tuned to address the variation in the parameters.
@@
In ${\mathbf{D}(\mathsf{v})}^{-1}-\mathbf{N}^{\mkern-1.5mu\mathsf{T}}{\mathbf{L}}_{0}^{-1}\mathbf{M}$, only the diagonal blocks ${D(\mathsf{v})}^{-1}\otimes I-\mathbf{N}_{1}^{\mkern-1.5mu\mathsf{T}}{\mathbf{L}}_{0}^{-1}\mathbf{M}_{1}$ and $I\otimes{D(\mathsf{v})}^{-1}-\mathbf{N}_{2}^{\mkern-1.5mu\mathsf{T}}{\mathbf{L}}_{0}^{-1}\mathbf{M}_{2}$ depend on the current parameters. Moreover, both ${D(\mathsf{v})}^{-1}\otimes I$ and $I\otimes{D(\mathsf{v})}^{-1}$ are diagonal matrices so that the sparsity pattern of $\mathbf{N}_{1}^{\mkern-1.5mu\mathsf{T}}{\mathbf{L}}_{0}^{-1}\mathbf{M}_{1}$ and $\mathbf{N}_{2}^{\mkern-1.5mu\mathsf{T}}{\mathbf{L}}_{0}^{-1}\mathbf{M}_{2}$ is preserved. In particular, ${D(\mathsf{v})}^{-1}\otimes I-\mathbf{N}_{1}^{\mkern-1.5mu\mathsf{T}}{\mathbf{L}}_{0}^{-1}\mathbf{M}_{1}$ is still a $k\times k$ block matrix with diagonal blocks whereas $I\otimes{D(\mathsf{v})}^{-1}-\mathbf{N}_{2}^{\mkern-1.5mu\mathsf{T}}{\mathbf{L}}_{0}^{-1}\mathbf{M}_{2}$ is block diagonal with $k\times k$ blocks on the diagonal. By exploiting such a significant sparsity pattern, we are thus able to efficiently solve linear systems with ${D(\mathsf{v})}^{-1}\otimes I-\mathbf{N}_{1}^{\mkern-1.5mu\mathsf{T}}{\mathbf{L}}_{0}^{-1}\mathbf{M}_{1}$ and $I\otimes{D(\mathsf{v})}^{-1}-\mathbf{N}_{2}^{\mkern-1.5mu\mathsf{T}}{\mathbf{L}}_{0}^{-1}\mathbf{M}_{2}$ by a sparse direct method, regardless of the change in $\mathsf{v}$.
@@
To conclude, the solution $X(\mathsf{v})$ to (1), in case of moderate $n$, can be computed as follows
@@
In Table 1 we summarize the operations that can be performed offline and online. Notice that the $\mathcal{O}(n^{3})$ flops needed to compute $W(\mathsf{v})$ come from the final matrix-matrix multiplications by $Q_{0}$ and $Q_{0}^{\mkern-1.5mu\mathsf{T}}$ in (16). These operations can be avoided for certain $f$, thus reducing the cost of computing $W(\mathsf{v})$ to $\mathcal{O}(n^{2})$ flops.

### v2

@@ Abstract
Sequences of parametrized Lyapunov equations can be encountered in many application settings. Moreover, solutions of such equations are often intermediate steps of an overall procedure whose main goal is the computation of $\operatorname*{trace}(EX)$ where $X$ denotes the solution of a Lyapunov equation and $E$ is a given matrix. We are interested in addressing problems where the parameter dependency of the coefficient matrix is encoded as a low-rank modification to a seed, fixed matrix. We propose two novel numerical procedures that fully exploit such a common structure. The first one builds upon the Sherman-Morrison-Woodbury (SMW) formula and recycling Krylov techniques, and it is well-suited for small dimensional problems as it makes use of dense numerical linear algebra tools. The second algorithm can instead address large-scale problems by relying on state-of-the-art projection techniques based on the extended Krylov subspace. We test the new algorithms on several problems arising in the study of damped vibrational systems and the analyses of output synchronization problems for multi-agent systems. Our results show that the algorithms we propose are superior to state-of-the-art techniques as they are able to remarkably speed up the computation of accurate solutions.
@@
Novel, efficient methods for solving sequences of parametrized Lyapunov equations have been designed. By taking full advantage of the parameter dependency structure, we are able to tackle both small-scale and large-scale problems. The numerical results show the superiority of the proposed methods when compared to state-of-the-art routines, especially in terms of computational time.
@@ 1 Introduction and setting
Here $A_{0},Q\in{\mathbb{R}}^{n\times n}$, $B_{l},B_{r}\in{\mathbb{R}}^{n\times k}$ are fixed matrices and the matrix $D(\mathsf{v})=\operatorname{diag}{(v_{1},v_{2},\ldots,v_{k})}$ contains parameters $v_{i}\in{\mathbb{R}}\setminus\left\{0\right\}$, for $i=1\ldots,k$, encoded in the parameter vector $\mathsf{v}=\begin{bmatrix}v_{1},v_{2},\ldots,v_{k}\end{bmatrix}^{\mkern-1.5mu\mathsf{T}}\in{({\mathbb{R}}\setminus\left\{0\right\})}^{k}$ [^1] For the sake of simplicity, we assume the data of our problem to be real. Our algorithms can handle complex data by applying straightforward modifications.. Note that, without loss of generality, we assume that the parameters in $\mathsf{v}$ are nonzero. The number of parameters $k$ is supposed to be very small, $k=\mathcal{O}(1)$, as this is the case in the application settings we are interested in; see Section 3.
@@
We are especially interested in devising efficient procedures for calculating $\operatorname*{trace}(EX(\mathsf{v}))$, for some matrix $E$, for many different parameter vectors $\mathsf{v}$. Indeed, $\operatorname*{trace}(EX(\mathsf{v}))$ often corresponds to the target value we are interested in computing, with calculating $X(\mathsf{v})$ only being an intermediate step. For instance this is the case when computing the $H_{2}$-norm of a linear system. An efficient computation of $\operatorname*{trace}(EX(\mathsf{v}))$ is crucial for obtaining competitive numerical schemes in terms of both computational performance and memory requirements for the applications mentioned above. Our method also covers efficient calculation of quantities of the form $f(X(\mathsf{v}))$, for a given function $f$ that is assumed to be additive, i.e. $f(X+Y)=f(X)+f(Y)$. To simplify the notation, in the sequel we will denote $\operatorname*{trace}(EX)$ by $f(X)$.
@@
The first procedure we suggest in this paper builds upon the work in [57, 31], which we enhance with a recycling Krylov technique. This routine makes use of dense linear algebra tools so that it is well suited for small dimensional problems, i.e., for moderate values of $n$. The second routine addresses the large-scale case employing projection methods for linear matrix equations.
@@
In Section 2 we present the main contribution of this work. In particular, a novel scheme combining the SMW formula with recycling Krylov-like methods is illustrated in Section 2.1. As already mentioned, this scheme is able to efficiently deal with small dimensional problems only as it employs dense numerical linear algebra tools like, e.g., the full eigenvalue decomposition of $A_{0}$, so that it cannot be applied in the large-scale setting. We address the latter scenario in Section 2.2, where a sophisticated projection technique, based on the extended Krylov subspace (17), is illustrated. In this framework, we will assume to be able to efficiently solve linear systems with $A_{0}$. Some details about two application settings where equations of the form (1) are encountered are reported in Section 3. In particular, damped vibrational systems are treated in Section 3.1, whereas in Section 3.2 multi-agent systems are considered. A panel of diverse numerical experiments displaying the effectiveness of our methodology is presented in Section 4, while Section 5 collects our conclusions.
@@ 2 The novel solution methods
2.1 Small-scale setting: SMW and recycling Krylov methods
@@
The strategy proposed in this section builds upon the work in [57, 31]. We extend their approach to the more general class of problems of the form (1), and we allow for more than one parameter, i.e., $k$ can be larger than one. The scheme we are going to present relies on the computation of the full eigenvalue decomposition of $A_{0}$, namely $A_{0}=Q_{0}\Lambda_{0}Q_{0}^{-1}$ with $Q_{0}\in{\mathbb{C}}^{n\times n}$ and $\Lambda_{0}=\operatorname{diag}(\lambda_{1},\ldots,\lambda_{n})$, $\lambda_{i}\in{\mathbb{C}}$, $i=1,\ldots,n$.
@@
The use of the Sherman-Morrison-Woodbury formula in the matrix setting is not new; see, e.g., [21, 32][12, Section 3][34, Section 5][4, 45]. However, applying the general-purpose schemes presented in these papers to our parameter-dependent framework would not lead to an efficient solution process. Indeed, this would be equivalent to explicitly computing (2.1) every time $\mathsf{v}$ changes. Our main goal here is to divide the operations that depend on $\mathsf{v}$ from those that do not, as much as possible, such that the $\mathsf{v}$-independent calculations can be performed offline, only once. In the interest of the reader, we thus spell out the details of our procedure, in spite of some similarities with prior work.
@@
While ${\mathbf{D}(\mathsf{v})}^{-1}=\begin{bmatrix}{D(\mathsf{v})}^{-1}\otimes I&\\ &I\otimes{D(\mathsf{v})}^{-1}\\ \end{bmatrix}$ is a diagonal matrix, the structure of $\mathbf{N}^{\mkern-1.5mu\mathsf{T}}{\mathbf{L}}_{0}^{-1}\mathbf{M}$ is more involved. A first study of its structure can be found in [57]. Unlike what is done in [57], here we exploit the diagonal pattern of ${\mathbf{L}}_{0}^{-1}$, which leads to a useful representation of the blocks of $\mathbf{N}^{\mkern-1.5mu\mathsf{T}}{\mathbf{L}}_{0}^{-1}\mathbf{M}$ when the latter is viewed as a $2\times 2$ block matrix. Such a block structure will help us to design efficient preconditioners for the solution of the linear system in (10) as well.
@@
where $\delta_{q,s}$ denotes the Kronecker delta. The relation in (2.1) shows that $\mathbf{N}_{1}^{\mkern-1.5mu\mathsf{T}}{\mathbf{L}}_{0}^{-1}\mathbf{M}_{1}$ is a $k\times k$ block matrix whose blocks are all diagonal. In particular, the block in the $(t,p)$ position is given by the matrix $\mathop{\mathrm{diag}}\left({\mathcal{L}}\left((Q_{0}^{\mkern-1.5mu\mathsf{T}}B_{r}e_{t})\mathop{\circ}(Q_{0}^{-1}B_{l}e_{p})\right)\right)$.
@@
Unlike its diagonal blocks, the off-diagonal blocks of $\mathbf{N}^{\mkern-1.5mu\mathsf{T}}{\mathbf{L}}_{0}^{-1}\mathbf{M}$, namely $\mathbf{N}_{2}^{\mkern-1.5mu\mathsf{T}}{\mathbf{L}}_{0}^{-1}\mathbf{M}_{1}$, and $\mathbf{N}_{1}^{\mkern-1.5mu\mathsf{T}}{\mathbf{L}}_{0}^{-1}\mathbf{M}_{2}$, do not possess a structured sparsity pattern. Nevertheless, we can still apply the same strategy we employed above to construct $\mathbf{N}_{2}^{\mkern-1.5mu\mathsf{T}}{\mathbf{L}}_{0}^{-1}\mathbf{M}_{1}$ and $\mathbf{N}_{1}^{\mkern-1.5mu\mathsf{T}}{\mathbf{L}}_{0}^{-1}\mathbf{M}_{2}$ while avoiding the explicit computation of $\mathbf{M}_{1}$, $\mathbf{M}_{2}$, $\mathbf{N}_{1}$, and $\mathbf{N}_{2}$. If $i=(s-1)k+t$ and $j=(p-1)n+q$, we have
@@
We employ the $s_{\ell}$ eigenvectors corresponding to the $s_{\ell}$ eigenvalues of smallest magnitude [^2] This is the default strategy in [43]., but different approximate eigenvectors can be used as well; see [42, Section 2.4].
The convergence of GCRO-DR is driven by the spectral properties of ${\mathbf{D}(\mathsf{v})}^{-1}-\mathbf{N}^{\mkern-1.5mu\mathsf{T}}{\mathbf{L}}_{0}^{-1}\mathbf{M}$. However, the latter ones are extremely tricky to identify in general. Only by allowing further assumptions on $B_{l}$, $B_{r}$, and $A_{0}$, we may come up with sensible statements. For instance, if $B_{l}=B_{r}$ and $A_{0}$ is Hermitian, ${\mathbf{D}(\mathsf{v})}^{-1}-\mathbf{N}^{\mkern-1.5mu\mathsf{T}}{\mathbf{L}}_{0}^{-1}\mathbf{M}$ is Hermitian too and GCRO-DR can take advantage of that; see [42, Theorem 3.1]. Similarly, if, in addition, all the parameters $v_{i}$’s are, e.g., strictly positive and $A_{0}$ is stable, then ${\mathbf{D}(\mathsf{v})}^{-1}-\mathbf{N}^{\mkern-1.5mu\mathsf{T}}{\mathbf{L}}_{0}^{-1}\mathbf{M}$ is also positive definite. GCRO-DR can take advantage of this property as well.
GCRO-DR allows for preconditioning. In particular, thanks to the pattern of ${\mathbf{D}(\mathsf{v})}^{-1}-\mathbf{N}^{\mkern-1.5mu\mathsf{T}}{\mathbf{L}}_{0}^{-1}\mathbf{M}$ we are able to design efficient preconditioning operators, which can be cheaply tuned to address the variation in the parameters.
@@
In ${\mathbf{D}(\mathsf{v})}^{-1}-\mathbf{N}^{\mkern-1.5mu\mathsf{T}}{\mathbf{L}}_{0}^{-1}\mathbf{M}$, only the diagonal blocks ${D(\mathsf{v})}^{-1}\otimes I-\mathbf{N}_{1}^{\mkern-1.5mu\mathsf{T}}{\mathbf{L}}_{0}^{-1}\mathbf{M}_{1}$ and $I\otimes{D(\mathsf{v})}^{-1}-\mathbf{N}_{2}^{\mkern-1.5mu\mathsf{T}}{\mathbf{L}}_{0}^{-1}\mathbf{M}_{2}$ depend on the current parameters. Moreover, both ${D(\mathsf{v})}^{-1}\otimes I$ and $I\otimes{D(\mathsf{v})}^{-1}$ are diagonal matrices so that the sparsity pattern of $\mathbf{N}_{1}^{\mkern-1.5mu\mathsf{T}}{\mathbf{L}}_{0}^{-1}\mathbf{M}_{1}$ and $\mathbf{N}_{2}^{\mkern-1.5mu\mathsf{T}}{\mathbf{L}}_{0}^{-1}\mathbf{M}_{2}$ is preserved. In particular, ${D(\mathsf{v})}^{-1}\otimes I-\mathbf{N}_{1}^{\mkern-1.5mu\mathsf{T}}{\mathbf{L}}_{0}^{-1}\mathbf{M}_{1}$ is still a $k\times k$ block matrix with diagonal blocks whereas $I\otimes{D(\mathsf{v})}^{-1}-\mathbf{N}_{2}^{\mkern-1.5mu\mathsf{T}}{\mathbf{L}}_{0}^{-1}\mathbf{M}_{2}$ is block-diagonal with $k\times k$ blocks on the diagonal. By exploiting such a significant sparsity pattern, we are thus able to efficiently solve linear systems with ${D(\mathsf{v})}^{-1}\otimes I-\mathbf{N}_{1}^{\mkern-1.5mu\mathsf{T}}{\mathbf{L}}_{0}^{-1}\mathbf{M}_{1}$ and $I\otimes{D(\mathsf{v})}^{-1}-\mathbf{N}_{2}^{\mkern-1.5mu\mathsf{T}}{\mathbf{L}}_{0}^{-1}\mathbf{M}_{2}$ by a sparse direct method, regardless of the change in $\mathsf{v}$.
@@
The solution $X(\mathsf{v})$ to (1), in case of moderate $n$, can thus be computed as follows,
@@
In Table 1 we summarize the operations that can be performed offline and online. Notice that the $\mathcal{O}(n^{3})$ flops needed to compute $W(\mathsf{v})$ come from the final matrix-matrix multiplications by $Q_{0}$ and $Q_{0}^{\mkern-1.5mu\mathsf{T}}$ in (16). These operations can be avoided for $f=\operatorname*{trace}$ and $Q_{0}$ orthogonal, thus reducing the cost of computing $W(\mathsf{v})$ to $\mathcal{O}(n^{2})$ flops.
@@
We conclude this section by recalling the reader that $w(\mathsf{v})$ has been computed by GCRO-DR, an iterative method that has been run up to a certain tolerance on the relative residual norm associated to (14). Therefore, we can think of $w(\mathsf{v})$ as being of the form $w(\mathsf{v})=w_{*}(\mathsf{v})+\ell(\mathsf{v})$, where $w_{*}(\mathsf{v})$ is the exact solution to (14) whereas $\ell(\mathsf{v})$ is the error vector coming from the computation of $w(\mathsf{v})$. The magnitude of $\ell(\mathsf{v})$ plays an important role in the overall accuracy that can be attained by $X(\mathsf{v})$ in (16). In particular, by writing $\ell(\mathsf{v})=[\ell_{1}(\mathsf{v});\ell_{2}(\mathsf{v})]$, $\ell_{1}(\mathsf{v}),\ell_{2}(\mathsf{v})\in\mathbb{C}^{nk}$, and $E_{1}(\mathsf{v})\in\mathbb{C}^{n\times k}$, $\ell_{1}(\mathsf{v})=\text{vec}(E_{1}(\mathsf{v}))$, $E_{2}(\mathsf{v})\in\mathbb{C}^{k\times n}$, $\ell_{2}(\mathsf{v})=\text{vec}(E_{2}(\mathsf{v}))$, a direct computation shows that we can write
$X(\mathsf{v})=X_{0}+Z(\mathsf{v})+W_{*}(\mathsf{v})+E(\mathsf{v}),$
where $E(\mathsf{v})=Q_{0}\left({\mathcal{L}}\mathop{\circ}\left(E_{1}(\mathsf{v})B_{l}^{\mkern-1.5mu\mathsf{T}}Q_{0}^{-\mkern-1.5mu\mathsf{T}}+Q_{0}^{-1}B_{l}E_{2}(\mathsf{v})^{\mkern-1.5mu\mathsf{T}}\right)\right)Q_{0}^{\mkern-1.5mu\mathsf{T}}$. To bound $\|E(\mathsf{v})\|_{F}$ it may be more convenient to look at $E(\mathsf{v})$ as the solution to the following Lyappunov equation
$A_{0}E(\mathsf{v})+E(\mathsf{v})A_{0}^{\mkern-1.5mu\mathsf{T}}=Q_{0}E_{1}(\mathsf{v})B_{l}^{\mkern-1.5mu\mathsf{T}}+B_{l}E_{2}(\mathsf{v})^{\mkern-1.5mu\mathsf{T}}Q_{0}^{\mkern-1.5mu\mathsf{T}}.$
Indeed, for stable $A_{0}$, by defining $\alpha_{0}$ as the largest eigenvalue of the symmetric part of $A_{0}$, namely $\alpha_{0}=\lambda_{\max}((A_{0}+A_{0}^{\mkern-1.5mu\mathsf{T}})/2)$, if $\alpha_{0}<0$, standard results on the norm of solutions to Lyapunov equations say that
$\|E(\mathsf{v})\|_{F}\leqslant\frac{1}{2|\alpha_{0}|}\|Q_{0}E_{1}(\mathsf{v})B_{l}^{\mkern-1.5mu\mathsf{T}}+B_{l}E_{2}(\mathsf{v})^{\mkern-1.5mu\mathsf{T}}Q_{0}^{\mkern-1.5mu\mathsf{T}}\|_{F};$
see, e.g., [62].